The cost of the gold bracelet was20.00 dollars more than three times the silver one.Together they cost 320.00dollars. How much did each cost?
Answer by robertb(5830)

silver = s
gold = 3s + 20. also, s+3s+20 = 320. Then
4s + 20 = 320,
4s = 300,
s = 75, the cost of the silver bracelet.
3s + 20 = 245, the cost of the gold bracelet.